Job Costs Using a Plantwide Overhead Rate

Naranjo Company designs industrial prototypes for outside companies. Budgeted overhead for the year was $360,000, and budgeted direct labor hours were 24,000. The average wage rate for direct labor is expected to be $30 per hour. During June, Naranjo Company worked on four jobs. Data relating to these four jobs follow:

Job 39 Job 40 Job 41 Job 42
Beginning balance $25,500 $32,900 $16,800 $1,400
Materials requisitioned 19,400 23,400 8,700 15,700
Direct labor cost 10,500 20,500 3,350 6,600

Overhead is assigned as a percentage of direct labor cost. During June, Jobs 39 and 40 were completed; Job 39 was sold at 130 percent of cost. (Naranjo had originally developed Job 40 to order for a customer; however, that customer was near bankruptcy and the chance of Naranjo being paid was growing dimmer. Naranjo decided to hold Job 40 in inventory while the customer worked out its financial difficulties. Job 40 is the only job in Finished Goods Inventory.) Jobs 41 and 42 remain unfinished at the end of the month.

Required:

1. Calculate the overhead rate based on direct labor cost.

% of direct labor cost


Hide

2. Set up a simple job-order cost sheet for all jobs in process during June.

Naranjo Company

Job-Order Cost Sheets

   

Job 39

Job 40

Job 41

Job 42

Balance, June 1

$  

$  

$  

$  

  

  

  

  

  

  

  

  

  

  

  

  

Total

$  

$  

$  

$  

3. What if the expected direct labor rate at the beginning of the year was $24 instead of $30? What would the overhead rate be? If required, round your answer to one decimal place.

% of direct labor cost

How would the cost of the jobs be affected?

- Select your answer -Applied overhead would increaseApplied overhead would decreaseApplied overhead would not changeCorrect 2 of Item 3
Answer #1
1] Estimated direct labor cost = 24000*30 = $     7,20,000
Budgeted overhead $     3,60,000
Overhead rate = 360000/720000 = 50.00% of Direct labor cost
2] Job 39 Job 40 Job 41 Job 42
Balance, June 1 $    25,500 $ 32,900 $   16,800 $     1,400
Direct Material cost $      9,400 $ 23,400 $     8,700 $   15,700
Direct labor cost $    10,500 $ 20,500 $     3,350 $     6,600
Overhead assigned $           5,250 $      10,250 $         1,675 $          3,300
Total $         50,650 $      87,050 $       30,525 $       27,000
3] Estimated direct labor cost = 24000*24 = $     5,76,000
Budgeted overhead $     3,60,000
Overhead rate = 360000/576000 = 62.50% of Direct labor cost
How would the cost of jobs be affected?
Applied overhead would not change.
The reason is that, the direct labor cost will
get reduced to the extent that the new
overhead rate when applied to the revised
cost of the jobs will yield the same OH
assignment for the jobs.